Problem # 3. A pair of packets is sent over the network to an end user. It is known that the first packet will be lost with probability 1/3. If the first packet is received, then the second packet is lost with probability 1/6. If the first packet is lost then the second packet is lost with probability 2/3. Find the probabilities of the events:

a) Exactly one packet is received by the end user.

b) Both packets are lost.
Answer #1

Tree Diagram

a) P(exactly one packet is received by end user) = P(first is lost and second is received) + P(first is received and second is lost)

= 1/3 x 1/3 + 2/3 x 1/6

= 2/9

b) P(both packets are lost) = 1/3 x 2/3

= 2/9